OpenStudy (anonymous):

Medal & fan. Determine the solution set of (x + 2)^2 = 36. Would i use foil here?

OpenStudy (freckles):

Take square root of both sides.

OpenStudy (freckles):

you will wind up with two equations to solve x+2=6 or x+2=-6

OpenStudy (anonymous):

x=4 or x=-4?

OpenStudy (hybrik):

Yes, x = + or - 4 [Also known as -4,4]

OpenStudy (anonymous):

answer choices are: a.) (6,-8) b.) (4,-8) c.) (4,-6)

OpenStudy (hybrik):

What is the square root of 36?

OpenStudy (anonymous):

\[\pm6\]

OpenStudy (anonymous):

in this case negative

OpenStudy (hybrik):

So find what x+2 = 6 and x-2 = 6, solve and you get, x=4, and x=8

OpenStudy (anonymous):

its for sure positive 8?

OpenStudy (hybrik):

4,-8

OpenStudy (anonymous):

oh okay

OpenStudy (freckles):

\[x+2=6 \text{ or } x+2=-6 \\ \text{ subtracting 2 on both sides gives } \\ x=6-2 \text{ or } x=-6-2 \\ x=4 \text{ or } x=-8\]

OpenStudy (freckles):

not 8

OpenStudy (freckles):

pluggin in 8 yields (8+2)^2=36 which is not a true equation

OpenStudy (freckles):

(4+2)^2=36 is true (-8+2)^2=36 is also true @yungjvke